# Unleashing the Power of Task-Specific Directions in Parameter Efficient Fine-tuning

### v3

@@ D.1 Implementation Details
These insights affirm the utility of $\Delta\mathbf{W}$ in accurately capturing and representing the information of real TSDs. The method shows potential in leveraging learned directional changes to identify and utilize key features specific to given tasks, underscoring the robustness of $\Delta\mathbf{W}$ as a predictive tool in practical applications.
For implementation details, please refer to Sec. D.4.
Table 6: The template of each dataset used to create commonsense reasoning data for fine-tuning.
Dataset Fine-tuning Data Template
BoolQ Please answer the following question with true or false, question: [QUESTION]
Answer format: true/false
the correct answer is [ANSWER]
PIQA Please choose the correct solution to the question: [QUESTION]
Solution1: [SOLUTION_1]
Solution2: [SOLUTION_2]
Answer format: solution1/solution2
the correct answer is [ANSWER]
SIQA Please choose the correct solution to the question: [QUESTION]
Answer1: [ANSWER_1]
Answer2: [ANSWER_2]
Answer3: [ANSWER_3]
Answer format: answer1/answer2/answer3
the correct answer is [ANSWER]
HellaSwag Please choose the correct ending to complete the given sentence: [ACTIVITY_LABEL]: [CONTEXT]
Ending1: [ENDING_1]
Ending2: [ENDING_2]
Ending3: [ENDING_3]
Ending4: [ENDING_4]
Answer format: ending1/ending2/ending3/ending4
the correct answer is [ANSWER]
WinoGrande Please choose the correct answer to fill in the blank to complete the given sentence: [SENTENCE]
Option1: [OPTION_1]
Option2: [OPTION_2]
the correct answer is [ANSWER]
ARC-e & ARC-c Please choose the correct solution to the question: [QUESTION]
Answer1: [ANSWER_1]
Answer2: [ANSWER_2]
Answer3: [ANSWER_3]
Answer4: [ANSWER_4]
Answer format: answer1/answer2/answer3/answer4
the correct answer is [ANSWER]
OBQA Please choose the correct solution to the question: [QUESTION]
Answer1: [ANSWER_1]
Answer2: [ANSWER_2]
Answer3: [ANSWER_3]
Answer4: [ANSWER_4]
Answer format: answer1/answer2/answer3/answer4
the correct answer is [ANSWER]
Table 7: Hyper-parameter settings of LoRA-Dash on commonsense reasoning task.
Hyper-parameters LLaMA-7B LLaMA2-7B LLaMA3-8B
Rank $r$ 4 8 16 32 64 16 32 16 32
$\alpha$ 8 16 32 64 128 32 64 32 64
LR 5e-4 4e-4 5e-4 1e-4 0.9e-4 2e-4 1e-4 2e-4 0.8e-4
LR Scheduler Linear
Dropout 0.05
Optimizer AdamW
Batch size 16
Warmup Steps 100
Epochs 3
Where Q, K, V, Up, Down
Table 8: An example from BoolQ. The outputs of LLaMA-7B with LoRA and LoRA-Dash.
              Question               was whiskey galore based on a true story?
              Label               true
              LoRA               the correct answer is false
              LoRA-Dash               the correct answer is true
Table 9: An example from PIQA. The outputs of LLaMA-7B with LoRA and LoRA-Dash.
Question To add chocolate and caramel to hot milk and cream.
Solution1: Chop the chocolate into small chunks.
Bring the milk and cream to a simmer on a low-medium heat.
Once the mixture begins to cool, remove from the heat and add the chocolate and caramel sauce.
Solution2: Chop the chocolate into small chunks.
Bring the milk and cream to a simmer on a low - medium heat.
Once the mixture begins to simmer, remove from the heat and add the chocolate and caramel sauce.
Label solution2
LoRA the correct answer is solution1
LoRA-Dash the correct answer is solution2
Table 10: An example from SIQA. The outputs of LLaMA-7B with LoRA and LoRA-Dash.
Question Lee challenged her to a fight after she talked bad about Lee to his friends.
How would Others feel as a result?
Answer1: happy that Lee wants a fight
Answer2: angry
Answer3: fight
Label answer2
LoRA the correct answer is answer1
LoRA-Dash the correct answer is answer2
Table 11: An example from HellaSwag. The outputs of LLaMA-7B with LoRA and LoRA-Dash.
Question Personal Care and Style: [header] How to dress rugged (men) [title] Shop for pants.
[step] Great rugged outfits start with the right pants.
Shop for jeans, khakis and canvas with straight-cut legs, and avoid any flashy graphics or loud colors.
Ending1: Also, make sure you try on pants to ensure they fit comfortably.
Rugged styles are about utility, so the pants shouldn’t be extremely tight or extremely baggy:
look for “relaxed fit” or “straight-cut”.
Ending2: Men should try slacks, dress pants, or street clothes for a rugged look.
[substeps] Choose gut-length, flared, or boot-cut trousers.
Ending3: A good rule of thumb is to dress in the right pair of pants, and balance it with that.
Consider clothing tucked into your upper thighs with high-rise, thigh high, or flare pants.
Ending4: Lightweight and comfortable feet are always a major consideration when shopping for a rugged shirt.
[substeps] These hiking boots, while lightweight and appropriate for longer, are a must.
Label ending1
LoRA the correct answer is ending2
LoRA-Dash the correct answer is ending1
Table 12: An example from WinoGrande. The outputs of LLaMA-7B with LoRA and LoRA-Dash.
Question Lawrence was always gaining weight while Dennis was losing it as _ liked to eat too little.
Option1: Lawrence
Option2: Dennis
Label option2
LoRA the correct answer is option1
LoRA-Dash the correct answer is option2
Table 13: An example from ARC-e. The outputs of LLaMA-7B with LoRA and LoRA-Dash.
Question There are a variety of events that can disrupt ecosystems.
Some events result in local disturbances and others can affect a much larger area.
Which event would most likely result in change on a global scale?
Answer1: land scarred by forest fires
Answer2: erosion of ocean shorelines
Answer3: converging plate boundaries
Answer4: destruction of tropical rain forests
Label answer4
LoRA the correct answer is answer3
LoRA-Dash the correct answer is answer4
Table 14: An example from ARC-c. The outputs of LLaMA-7B with LoRA and LoRA-Dash.
         Question          What is a similarity between sound waves and light waves?
         Answer1: Both carry energy.
         Answer2: Both travel in vacuums.
         Answer3: Both are caused by vibrations.
         Answer4: Both are traveling at the same speed.
         Label          answer1
         LoRA          the correct answer is answer3
         LoRA-Dash          the correct answer is answer1
Table 15: An example from OBQA. The outputs of LLaMA-7B with LoRA and LoRA-Dash.
Question Which best demonstrates the concept of force causing an increase in speed?
Answer1: skating on a rough surface
Answer2: a full bag swung in circles
Answer3: a computer powering on
Answer4: a baker stirring batter
Label answer2
LoRA the correct answer is answer1
LoRA-Dash the correct answer is answer2
D.4 Details on Commonsense Reasoning Task
The commonsense reasoning benchmarks consist of 8 distinct sub-tasks, each with its designated dataset, i.e., BoolQ Clark et al. (2019), PIQA Bisk et al. (2020), SIQA Sap et al. (2019), HellaS. Zellers et al. (2019), WinoG. Sakaguchi et al. (2021), ARC-e/ARC-c Clark et al. (2018), OBQA Mihaylov et al. (2018). Adhering to the protocol outlined by Hu et al. (2023), we merge the training datasets from all tasks to form a comprehensive training dataset (Commonsense170K dataset), subsequently performing evaluations against each individual task’s testing set. The template to create commonsense reasoning task is shown in Table. 6.
We fine-tune LLaMA-7B Touvron et al. (2023a), LLaMA2-7B Touvron et al. (2023b) and LLaMA3-8B AI@Meta (2024) on this task. Additionally, we integrate results from ChatGPT’s implementation with the gpt-3.5-turbo API, particularly focusing on zero-shot Chain of Thought approaches Wei et al. (2022).
To ensure equitable comparisons, initial fine-tuning for models utilizing LoRA-Dash is conducted under the LoRA configurations while only varying the learning rate to optimize performance. The hyper-parameter settings of LoRA-Dash are shown in Table. 7. The results of LoRA are cited from Hu et al. (2023); Liu et al. (2024).
We also represent some answer samples with fine-tuned LLaMA-7B. The answers are shown in Tables. 8-15.
Table 16: Details of GLUE dataset.
Dataset Task # Train # Dev # Test # Label Metrics
Single-Sentence Classification
CoLA Acceptability 8.5k 1k 1k 2 Matthews corr
SST-2 Sentiment 67k 872 1.8k 2 Accuracy
Similarity and Paraphrase
MRPC Paraphrase 3.7k 408 1.7k 2 Accuracy / F1
QQP Paraphrase 364k 40k 391k 2 Accuracy / F1
STS-B Similarity 7k 1.5k 1.4k 1 Pearson/ Spearman Corr
Natural Language Inference
MNLI NLI 393k 20k 20k 3 Accuracy
QNLI QA/NLI 108k 5.7k 5.7k 2 Accuracy
RTE NLI 2.5k 276 3k 2 Accuracy
Table 17: Hyper-parameter settings of LoRA-Dash on NLU task.
Hyper-parameter MNLI SST-2 CoLA QQP QNLI RTE MRPC STS-B
Optimizer AdamW
Warmup Ratio 0.1
LR schedule Linear
Rank $r$ 2 & 8
LoRA alpha 4 & 16
Max Seq. Len. 256 128 64 320 512 320 320 128
Batch Size 32 32 32 32 32 32 32 32
Learning Rate 5e-4 8e-4 8e-4 1e-3 5e-4 1.2e-3 1e-3 5e-4
Epochs 12 24 25 5 5 50 30 25
D.5 Details on Natural Language Understanding Task
For natural language understanding (NLU) task, we adopt the General Language Understanding Evaluation (GLUE) Wang et al. (2018) benchmark, which is designed to test capabilities across various tasks. This benchmark consists of two single-sentence classification tasks, CoLA Warstadt et al. (2019) and SST-2 Socher et al. (2013), three similarity and paraphrase tasks, MRPC Dolan & Brockett (2005), QQP Wang et al. (2018), and STS-B Cer et al. (2017), and three natural language inference tasks, MNLI Williams et al. (2017), QNLI Rajpurkar et al. (2016), and RTE Dagan et al. (2005); Bar-Haim et al. (2006); Giampiccolo et al. (2007); Bentivogli et al. (2009). The details of these datasets are shown in Table. 16.
We fine-tune DeBERTaV3-base and DeBERTaV3-large He et al. (2021b) models on this task. The hyper-parameter settings for this task is shown in Table. 17.
D.6 Details on Subject-driven Generation Task
In our experiment, we fine-tune the text-to-image diffusion models specifically tailored for subject-driven generation tasks, as outlined in recent research Ruiz et al. (2023). The objective of this task is to generate images that adhere closely to prompts associated with a particular subject, defined by a few exemplar images. This involves initially fine-tuning a text-to-image model using image-text pairs where the text contains a unique identifier (e.g., “A photo of a [V] cat”). Subsequent image generations are driven by new prompts incorporating this identifier, aiming to produce images aligned with the learned subject.
For this experiment, we use the SDXL5 model Podell et al. (2023), applying both LoRA and LoRA-Dash methods for fine-tuning. The fine-tuning process is conducted with a learning rate of 1e-4 and a batch size of 4. We train the model over 500 steps on a single 80GB A100 GPU, taking approximately 23 minutes to complete. For the generation phase, we execute 50 inference steps for each given prompt to synthesize the final images, which takes approximately 7 seconds to complete.
We mainly adopt the official data of DreamBooth Ruiz et al. (2023) for diffusion.
D.7 Details on the Amplification of TSD in Sec. 6.2
To explore to what extent LoRA-Dash can amplify the information of the TSDs it determines, we project the pretrained weights $\mathbf{W}$ and the weights merged after training, i.e., $\mathbf{W}+\Delta\mathbf{W}_{all}$, onto the LTSDs. Let $\bar{\mathbf{U}}=\begin{bmatrix}\bar{\mathbf{u}}_{1}&\ldots&\bar{\mathbf{u}}_{s}\end{bmatrix}$ and $\bar{\mathbf{V}}=\begin{bmatrix}\bar{\mathbf{v}}_{1}&\ldots&\bar{\mathbf{v}}_{s}\end{bmatrix}$, we can then define the amplification factor of $\Delta\mathbf{W}_{all}$ as the Frobenius norms of these two weights on these directions, which is given by:
$\frac{\|\bar{\mathbf{U}}^{\mathsf{T}}(\mathbf{W}+\Delta\mathbf{W}_{all})\bar{\mathbf{V}}\|_{F}}{\|\bar{\mathbf{U}}^{\mathsf{T}}\mathbf{W}\bar{\mathbf{V}}\|_{F}}.$ (9)
Specifically, we also evaluate the separate impacts of $\Delta\mathbf{W}_{AB}$ and $\Delta\mathbf{W}_{dash}$ in the amplification, with the corresponding amplification factor being
$\frac{\|\bar{\mathbf{U}}^{\mathsf{T}}(\mathbf{W}+\Delta\mathbf{W}_{AB})\bar{\mathbf{V}}\|_{F}}{\|\bar{\mathbf{U}}^{\mathsf{T}}\mathbf{W}\bar{\mathbf{V}}\|_{F}}\ \ {\rm and}\ \ \frac{\|\bar{\mathbf{U}}^{\mathsf{T}}(\mathbf{W}+\Delta\mathbf{W}_{dash})\bar{\mathbf{V}}\|_{F}}{\|\bar{\mathbf{U}}^{\mathsf{T}}\mathbf{W}\bar{\mathbf{V}}\|_{F}}.$ (10)
The results of the amplification are shown in Table. 4.
Moreover, as shown in Sec. B.2, TSDs are indeed the projection of $\Delta\mathbf{W}^{*}$’s direction onto the core bases of $\mathbf{W}$. Therefore, we also explored the amplification factor along the direction of $\Delta\mathbf{W}_{all}$. We replaced $\bar{\mathbf{U}}$ and $\bar{\mathbf{V}}$ in the previous amplification factor calculation with the top $r$ left and right singular vectors of $\Delta\mathbf{W}_{all}$, where $r$ is the rank setting of LoRA-Dash. We reported the corresponding results in Table. 18, where we can draw a conclusion:
Table 18: Amplification of $\Delta\mathbf{W}_{all}$ of LoRA-Dash onto TSDs or $\Delta\mathbf{W}_{all}$’s direction. We report the amplification factor from the self-attention key (k) projection of the first, middle and last layer of LLaMA-7B, as well as the average of all q, k, and v modules.
$\mathbf{U}$ and $\mathbf{V}$ from $r=8$ $r=16$ $r=32$
TSD $\Delta\mathbf{W}_{all}$ TSD $\Delta\mathbf{W}_{all}$ TSD $\Delta\mathbf{W}_{all}$
1st 282.15 22.75 628.60 30.13 42.00 3.07
16th 11.15 144.89 13.07 119.18 2.16 9.92
32nd 68.02 75.81 30.87 120.76 8.19 10.63
Avg. 22.85 153.49 32.12 126.26 5.28 11.31
  • The amplification factor on the directions of $\Delta\mathbf{W}_{all}$ is significantly higher than that on the TSDs. Since TSDs can be considered as a projection of $\Delta\mathbf{W}$ onto the coordinate system of $\mathbf{W}$, it is natural that some information is lost in this projection for we only select $s$ TSDs, leading to a lower amplification factor on the TSD compared to before the projection.
Additionally, we observed an interesting trend: the amplification factor on $\Delta\mathbf{W}_{all}$’s directions significantly decreases as the rank increases. In LoRA Hu et al. (2021), the authors claimed that the occurrence of a small number of TSDs leads to this situation. However, a more plausible explanation is that higher rank corresponds to $\Delta\mathbf{W}$ containing more irrelevant directions, thereby reducing the amplification factor.
To validate this, consider the setting of $r=32$ of $\Delta\mathbf{W}_{all}$. The 32 directions of $\Delta\mathbf{W}_{all}$ include some irrelevant directions, which is why the amplification factor is lower. This implies that if the irrelevant directions among these 32 are removed, the amplification factor could potentially increase. To explore whether this is the case, we recalculated the amplification factors using the top 2, 4, 8 and 16 singular directions of $\Delta\mathbf{W}_{all}$, and the results are shown in Table. 19. The observation that the amplification factor decreases as projections are made onto an increasing number of the top singular directions of $\Delta\mathbf{W}_{all}$ indicates that $\Delta\mathbf{W}_{all}$’s singular directions do indeed contain irrelevant directions in its bottom singular directions. It is not surprising, since TSDs are only a few directions, as shown in its properties. Therefore, it also motivates us to select several top directions in pre-launch phase of LoRA-Dash.
However, we believe that there is no necessary correlation between amplification factors and the effectiveness of a model. Amplification factors simply reflect how much the model amplifies features in certain directions, and the specific degree to which this should occur is unknown.
[…]

### v4

@@ 3.4 Challenges and Explorations: Using TSDs is Not as Easy as Expected
These insights affirm the utility of $\Delta\mathbf{W}$ in accurately capturing and representing the information of real TSDs. It shows potential in leveraging learned directional changes to identify and utilize key features specific to given tasks, underscoring the robustness of $\Delta\mathbf{W}$ as a predictive tool in practical applications. It leads to an important conclusion:
Observation 1.
This suggests that even without prior knowledge of TSDs, we can still capture their crucial information through the $\Delta\mathbf{W}$ obtained during LoRA training!
4 Unleash the Power of Task-specific Directions
To further unleash the power of TSDs in downstream tasks, this section introduces two methods utilizing TSDs. Both methods are divided into two phases: the “pre-launch phase”, where TSDs are identified, and the “dash phase”, which unleash the potential of theses identified TSDs. Changes occur during the dash phase. Specifically, the first method is LoRA-Dash, which directly simulates the changes in TSDs, while the other, called LoRA-Init, uses TSDs to initialize the matrices $\mathbf{A}$ and $\mathbf{B}$ in LoRA. The frameworks of LoRA-Dash and LoRA-Init are shown in Fig. 4.
Figure 4: Frameworks of LoRA-TSD, LoRA-Dash and LoRA-Init.
4.1 Pre-Launch Phase
4.2 Dash Phase for LoRA-Dash
To leverage the identified LTSDs, LoRA-Dash directly learns the changes in their coordinates, denoted as $\Delta{\sigma}_{i}$ for the $i$-th $\bar{\mathbf{u}}_{i}\bar{\mathbf{v}}_{i}^{\mathsf{T}}$, to further fine-tune the model for downstream tasks. $\Delta{\sigma}_{i}$ is initialized as zero. Mathematically, this adjustment is represented by $\sum_{i=1}^{s}\Delta{\sigma}_{i}\bar{\mathbf{u}}_{i}\bar{\mathbf{v}}_{i}^{\mathsf{T}}$. Ultimately, the updated weight matrix is given by:
$\mathbf{W}+\Delta\mathbf{W}_{AB}+\Delta\mathbf{W}_{dash}=\mathbf{W}+\mathbf{A}\mathbf{B}+\sum_{i=1}^{s}\Delta{\sigma}_{i}\bar{\mathbf{u}}_{i}\bar{\mathbf{v}}_{i}^{\mathsf{T}}.$ (5)
This equation forms the basis of LoRA-Dash, which aims to fully utilize the power of LTSDs derived from LoRA. During training, $\mathbf{A}$, $\mathbf{B}$, and $\Delta{\sigma}_{i}$ are continuously updated.
To further enhance efficiency, we transit to matrix computations for determining change rates. Let ${\mathbf{U}}=\begin{bmatrix}\mathbf{u}_{1}&\ldots&\mathbf{u}_{n}\end{bmatrix}$ and ${\mathbf{V}}=\begin{bmatrix}\mathbf{v}_{1}&\ldots&\mathbf{v}_{m}\end{bmatrix}$ represent the left and right singular vectors for the weight matrix $\mathbf{W}$, we have
${\mathbf{U}}^{\mathsf{T}}\Delta\mathbf{W}\mathbf{V}=\begin{bmatrix}\mathbf{u}_{1}^{\mathsf{T}}\\ \vdots\\ \mathbf{u}_{n}^{\mathsf{T}}\end{bmatrix}\Delta\mathbf{W}\begin{bmatrix}\mathbf{v}_{1}&\cdots&\mathbf{v}_{m}\end{bmatrix}$ (6)
$=\begin{bmatrix}\mathbf{u}_{1}^{\mathsf{T}}\Delta\mathbf{W}\mathbf{v}_{1}&\ldots&\mathbf{u}_{1}^{\mathsf{T}}\Delta\mathbf{W}\mathbf{v}_{m}\\ \vdots&\mathbf{u}_{i}^{\mathsf{T}}\Delta\mathbf{W}\mathbf{v}_{j}&\vdots\\ \mathbf{u}_{n}^{\mathsf{T}}\Delta\mathbf{W}\mathbf{v}_{1}&\ldots&\mathbf{u}_{n}^{\mathsf{T}}\Delta\mathbf{W}\mathbf{v}_{m}\end{bmatrix}.$
Therefore, we can derive the change rates $\mathbf{\delta}=\begin{bmatrix}\delta_{1}&\ldots&\delta_{n}\end{bmatrix}^{\mathsf{T}}$, which are the rectangle diagonal elements of the matrix $(\mathbf{U}^{\mathsf{T}}\Delta\mathbf{W}\mathbf{V})\odot(\mathbf{\Sigma}^{-1}+\epsilon\mathbf{I}_{n\times m})$.
TABLE I: Results on commonsense reasoning tasks. We fine-tune LLaMA-7B [4], LLaMA2-7B [6] and LLaMA3-8B [7] on this task.
ChatGPT - - 73.1 85.4 68.5 78.5 66.1 89.8 79.9 74.8 77.0
LLaMA-7B Fully FT 100 69.9 84.2 78.9 92.3 83.3 86.6 72.8 83.4 81.4
LoRAr=4 0.10 2.3 46.1 18.3 19.7 55.2 65.4 51.9 57.0 39.5
LoRA-Init 0.10 68.4 81.3 78.1 77.9 78.8 78.8 62.8 78.4 75.6
LoRA-Dash 0.10 65.2 79.9 78.3 82.8 77.1 78.6 65.4 78.4 75.7
LoRA-TSD 0.10 66.7 80.3 78.3 81.9 78.8 79.2 64.7 79.1 76.1
LoRAr=8 0.21 31.3 57.0 44.0 11.8 43.3 45.7 39.2 53.8 40.7
LoRA-Init 0.21 69.5 81.9 78.1 78.8 79.8 79.2 64.9 79.2 76.2
LoRA-Dash 0.21 69.8 81.1 77.3 85.1 81.1 77.2 64.1 79.6 76.9
LoRA-TSD 0.21 69.8 81.5 78.2 84.9 80.7 78.6 64.5 79.6 77.2
LoRAr=16 0.42 69.9 77.8 75.1 72.1 55.8 77.1 62.2 78.0 70.9
LoRA-Init 0.42 68.2 80.3 77.6 75.8 79.3 77.9 62.9 77.4 74.9
LoRA-Dash 0.42 66.9 80.2 77.8 78.8 79.2 78.0 61.9 77.4 75.0
LoRA-TSD 0.42 69.8 80.2 78.1 78.7 79.2 78.0 62.0 78.4 75.6
LoRAr=32 0.83 68.9 80.7 77.4 78.1 78.8 77.8 61.3 74.8 74.7
LoRA-Init 0.83 69.4 82.3 78.1 85.4 81.3 80.9 65.7 78.6 77.7
LoRA-Dash 0.83 69.9 82.8 78.6 84.9 81.6 82.3 66.5 80.8 78.4
LoRA-TSD 0.83 69.8 82.6 78.7 85.1 81.6 81.4 66.9 81.0 78.4
LoRAr=64 1.66 66.7 79.1 75.7 17.6 78.8 73.3 59.6 75.2 65.8
LoRA-Init 1.66 68.7 81.0 78.8 81.9 81.5 79.9 65.6 79.0 77.0
LoRA-Dash 1.66 69.6 79.5 76.0 82.8 75.8 81.5 64.7 81.0 76.4
LoRA-TSD 1.66 69.2 80.6 77.9 82.5 79.4 81.2 65.1 80.4 77.0
LLaMA2-7B Fully FT 100 72.2 84.9 80.9 93.1 84.7 87.5 74.2 85.1 82.8
LoRAr=16 0.41 72.4 82.5 79.1 88.5 80.6 82.3 67.1 78.9 79.3
LoRA-Init 0.41 71.7 81.6 79.5 89.5 81.9 82.9 67.9 79.6 79.3
LoRA-Dash 0.41 70.9 82.2 80.5 90.2 80.1 83.5 68.9 80.8 79.6
LoRA-TSD 0.41 71.3 81.2 80.1 89.9 81.7 83.3 68.4 80.9 79.6
LoRAr=32 0.82 69.8 79.9 79.5 83.6 82.6 79.8 64.7 81.0 77.6
LoRA-Init 0.82 69.8 83.0 78.0 87.7 84.1 83.1 68.3 82.0 79.5
LoRA-Dash 0.82 71.0 75.7 79.3 91.1 78.6 84.2 69.8 78.8 78.6
LoRA-TSD 0.82 70.6 81.9 79.4 91.1 83.9 84.0 69.8 82.3 80.4
LLaMA3-8B Fully FT 100 75.3 89.9 81.5 95.8 87.6 91.6 79.3 87.4 86.1
LoRAr=16 0.35 72.3 86.7 79.3 93.5 84.8 87.7 75.7 82.8 82.8
LoRA-Init 0.35 72.2 88.3 80.9 93.3 87.0 87.7 75.8 82.8 83.4
LoRA-Dash 0.35 74.8 88.0 80.6 95.2 85.6 89.0 77.4 84.8 84.4
LoRA-TSD 0.35 74.8 88.3 80.9 95.5 86.9 90.4 78.3 85.0 85.0
LoRAr=32 0.70 70.8 85.2 79.9 91.7 84.3 84.2 71.2 79.0 80.8
LoRA-Init 0.70 73.6 85.5 80.0 94.6 85.7 88.6 76.1 83.0 83.4
LoRA-Dash 0.70 75.3 88.5 80.2 95.7 86.8 90.7 80.2 85.6 85.4
LoRA-TSD 0.70 75.0 88.6 80.9 96.1 87.0 89.9 80.3 85.9 85.5
4.3 Dash Phase for LoRA-Init
In the Dash Phase, in addition to directly simulating the changes in TSDs, we can also use TSDs to initialize the matrices $\mathbf{A}$ and $\mathbf{B}$ in LoRA, which is a practical problem in LoRA series [22, 24, 25]. Specifically, we perform an SVD decomposition on the pre-trained weights $\mathbf{W}$, which can be split into the top $r$ TSD components and non-TSD components, i.e.,
$\mathbf{W}=\sum_{i=1}^{r}\bar{\sigma}_{i}\bar{\mathbf{u}}_{i}\bar{\mathbf{v}}_{i}^{\mathsf{T}}+\mathbf{W}_{res}=\bar{\mathbf{U}}\bar{\mathbf{\Sigma}}\bar{\mathbf{V}}^{\mathsf{T}}+\mathbf{W}_{res}.$ (7)
Here, $\bar{\sigma}_{i}$ represents the corresponding singular value of $\bar{\mathbf{u}}_{i}\bar{\mathbf{v}}_{i}^{\mathsf{T}}$, and $\mathbf{W}_{res}$ denotes the remaining components, which are frozen during training. We then use the most significant components that require the greatest changes to initialize the matrices $\mathbf{A}$ and $\mathbf{B}$ as
$\mathbf{A}=\bar{\mathbf{U}}\bar{\mathbf{\Sigma}}^{1/2},\mathbf{B}=\bar{\mathbf{\Sigma}}^{1/2}\bar{\mathbf{V}}^{\mathsf{T}}.$ (8)
Ultimately, the updated weight matrix is given by:
$\mathbf{W}\rightarrow\mathbf{W}_{res}+\mathbf{A}\mathbf{B}.$ (9)
This equation forms the basis of LoRA-Init. During training, only $\mathbf{A}$ and $\mathbf{B}$ are continuously updated.
4.4 LoRA-TSD: Unleashing the Power of TSD
We ultimately combine the two methods to create a fully optimized approach that fully harnesses the power of TSD: LoRA-TSD. Specifically, after identifying TSDs, LoRA-TSD both initializes the LoRA matrices $\mathbf{AB}$ using TSDs and directly models the changes in the TSD coordinates. The update formula for LoRA-TSD is as follows:
$\mathbf{W}_{tsd}=\mathbf{W}_{res}+\mathbf{A}\mathbf{B}+\sum_{i=1}^{s}\Delta{\sigma}_{i}\bar{\mathbf{u}}_{i}\bar{\mathbf{v}}_{i}^{\mathsf{T}},$ (10)
where $\mathbf{A}$, $\mathbf{B}$ and $\Delta{\sigma}_{i}$ are trained during fine-tuning. LoRA-TSD combines the advantages of both LoRA-Dash and LoRA-Init, fully unleashing the potential of TSD.
5 How Much Can TSD Boost the Performance of LoRA?
TABLE II: Results with DeBERTaV3 fine-tuned on GLUE development set. “FT” represents fully fine-tuning, and “Base’ and “Large” represent DeBERTaV3-base and DeBERTaV3-large, respectively.
Method Params(%) MNLI SST-2 CoLA QQP QNLI RTE MRPC STS-B All
Acc Acc Mcc Acc Acc Acc Acc Corr Avg.
Base(FT) 100% 89.90 95.63 69.19 91.87 94.03 83.75 90.20 91.60 88.27
LoRAr=2 0.18% 90.03 93.92 69.15 90.61 93.37 87.01 90.19 90.75 88.13
LoRA-Init 0.18% 89.34 95.76 71.45 91.58 94.32 89.17 91.42 91.85 89.36
LoRA-Dash 0.18% 90.14 95.42 72.41 91.65 94.36 89.89 91.67 91.64 89.65
LoRA-TSD 0.18% 90.26 96.10 72.41 91.65 94.77 90.03 91.91 91.89 89.88
LoRAr=8 0.72% 89.80 93.69 69.30 91.78 92.97 86.28 90.68 91.62 88.27
LoRA-Init 0.72% 90.31 95.64 71.05 91.39 93.78 87.19 90.93 91.53 88.98
LoRA-Dash 0.72% 90.55 96.99 70.78 92.39 94.22 88.09 91.18 91.91 89.52
LoRA-TSD 0.72% 90.37 96.70 72.41 93.44 94.77 90.10 91.91 92.07 90.22
Large(FT) 100% 91.81 96.93 75.27 93.01 96.02 92.68 92.20 92.98 91.36
LoRAr=2 0.20% 91.33 95.87 73.89 91.84 95.14 91.69 90.68 92.85 90.41
LoRA-Init 0.20% 91.59 96.10 75.34 92.67 95.48 92.06 91.67 92.96 90.98
LoRA-Dash 0.20% 91.65 96.11 76.11 92.61 95.52 92.78 92.18 93.05 91.25
LoRA-TSD 0.20% 91.65 96.28 76.34 92.89 95.77 92.93 92.03 92.89 91.35
LoRAr=8 0.80% 91.38 96.33 74.48 92.54 95.48 92.05 91.17 92.92 90.79
LoRA-Init 0.80% 91.30 96.44 76.34 92.49 95.75 92.42 91.67 92.73 91.14
LoRA-Dash 0.80% 91.18 96.45 76.88 92.81 95.85 93.51 91.91 92.86 91.43
LoRA-TSD 0.80% 91.27 96.51 76.94 92.99 95.78 93.51 92.18 93.05 91.53
To explore the performance gains of LoRA-Dash, LoRA-Init, and LoRA-TSD, both numerical and visual comparisons are employed. The numerical experiments focus on commonsense reasoning and natural language understanding tasks, assessing performance improvements of our methods in standard natural language processing metrics. Visual comparisons focuses on subject-driven generation task, involving using a text-to-image diffusion model to learn specific concepts (i.e., task-specific concept), providing a clear, qualitative and intuitive understanding of the impact of TSDs on performance. The detailed experimental settings are shown in Secs. 8.1 and 8.3-8.5, respectively.
5.1 Numerical Results
The numerical results, as shown in Tables. I-II, demonstrate that LoRA-Dash, LoRA-Init and LoRA-TSD significantly outperform LoRA across all tasks and models, with LoRA-TSD achieving the best performance. This distinction is so pronounced that further elaboration on performance gains might be unnecessary. Therefore, we instead focus on several key insights derived beyond mere performance gains.
Robustness to Parameter Budget: (When fine-tuning LLaMA-7B) LoRA shows enhanced performance with an increased parameter budget, indicating its dependence on larger rank sizes. Conversely, our methods maintain strong performance even under limited parameter conditions, which suggests that TSDs are crucial for maximizing fine-tuning efficiency. This highlights its effectiveness in harnessing the full potential of TSDs, even in constrained settings.
Surpassing Fully Fine-Tuning Performance: While LoRA typically falls short of the results achieved by fully fine-tuning, our methods in some cases outperform fully fine-tuning. This intriguing observation may provide new insights into optimizing training strategies for fine-tuning.
For comparisons with other methods, please refer to Tables. XI-XII.
5.2 Visual Results
Figure 5: Comparison of generated images from LoRA and LoRA-TSD on subject-driven generation task. Our method consistently aligns more closely with the subjects in the input images and adheres better to the given prompts than LoRA.
Figure 7: Ablation study on CoLA when fine-tuning DeBERTaV3-large of LoRA-Dash. a. Ablation study on the directions for dash. Beyond the top, random, and bottom directions, we also explored adjusting all directions and a combined adjustment of both bottom and top directions. b. Ablation study on the length of pre-launch phase $t$. c. Ablation study on the number of directions $s$ for dash.
TABLE III: Performance comparison between LoRA and LoRA-TSD on CLIP text-image alignment and image similarity metrics.
Metric (%) LoRA LoRA-TSD
Mean CLIP Text-Image Score 32.0 33.1
Mean Image Similarity 76.4 78.6
The visual comparisons, as shown in Fig. 5, demonstrate the superior fidelity of images generated by our method in aligning with the subjects of the input images compared to those by standard LoRA. For instance, images generated by LoRA of a dog and a teapot showed significant deviations from the input, whereas our method’s outputs maintained high consistency with the original images. Additionally, the effectiveness of TSDs is particularly evident in the response of our method to given prompts. TSDs effectively encapsulate the specific information of each prompt, enabling our methods to precisely grasp the semantic essence of prompts such as rainbow scarf or wheat field. This capability allows LoRA-TSD to adeptly render complex themes and details within the images.
We also conducted a user study using the DreamBooth dataset. We selected 10 different categories and generated corresponding images using five prompts for each category. We then invited 36 experts to evaluate the images based on subject similarity and prompt consistency, with each expert evaluating randomly sampled 25 pairs of images. During evaluation, the pairs of images were shuffled, and the experts were unaware of the source model for each image. Overall, the images LoRA-TSD generated achieved an 88.11% approval rating. This suggests that the images generated by LoRA-TSD were more favorably recognized by experts w.r.t. subject similarity and prompt consistency.
Moreover, we also evaluated the performance using the text-image score evaluated by CLIP [48] and the image similarity score. The CLIP text-image score measures the alignment between textual descriptions and the corresponding images, while the image similarity score quantifies the similarity between the original image and the generated image. The results are presented in Table III, where it is evident that LoRA-TSD outperforms LoRA on both metrics. The images generated by LoRA-TSD exhibit better text alignment and a stronger correspondence with the original images, demonstrating that LoRA-TSD achieves superior results compared to LoRA.
5.3 Ablation Study of LoRA-Dash
We first investigated the effects of TSDs during the pre-launch phase compared to other locations, such as selecting top, bottom or random core directions. The results when fine-tuning LLaMA3-8B and DeBERTaV3-large, as shown in Figs. 6(a) and 7(a), clearly indicate that selecting TSDs yields the best performance, with top selections performing the worst. This is not difficult to understand: directly dashing TSDs maximizes the ability of $\mathbf{W}$ to adapt to downstream tasks. Since TSDs are generally located towards the end of the spectrum, choosing directions from the bottom naturally results in better outcomes than selecting from the top.
[…]